OpenStudy (anonymous):

Which is the equivalent to the following expression? (3m+2mn-n2)+(m2+4mn-n2) A: 4mn2+6mn-2n2 B: 4m4-6mn-2n4 C: 3m2-6mn+2n2 D: 4m2+6mn+2n2 and why?

OpenStudy (anonymous):

A because if you take that equation and add it together you get the same answer as "A"

OpenStudy (anonymous):

In the second half of the equation, is m raised to the 2nd?

OpenStudy (anonymous):

all of the numbers following the letters in the equation are exponents

OpenStudy (anonymous):

Unless the question is typed incorrectly, none of those answers are correct. (3m+2mn-n^2)+(m^2+4mn-n^2) 3m+m^2=3m^2 2mn+4mn=6mn -n^2+-n^2=0 Which results in 3m^2+6mn as your final answer.